The Morris Corporation has $700,000 of debt outstanding, and it pays an interest rate of 10% annually. Morris's annual sales are $3.5 million, its average tax rate is 30%, and its net profit margin on sales is 3%. If the company does not maintain a TIE ratio of at least 4 to 1, its bank will refuse to renew the loan and bankruptcy will result. What is Morris's TIE ratio? Round intermediate calculations to two decimal places. Round your answer to two decimal places.
Answer #1

Times Interest Earned Ratio or TIE = Earnings before interest and Taxes/Annual Interest Expense

Net profit Margin = 3%

Net Income = 3%*3,500,000 = $105,000

Income before taxes = 105,000/(1-30%) = $150,000

Add: Interest Expense = 700,000*10% = $70,000

Earnings before Interest and Tax = $220,000

Times Interest Earned Ratio = 220,000/70,000

= 3.1428

i.e. 3.14 times

Hence, the bank will refuse to renew the loan